OpenStudy (anonymous):

What are the intercepts of the equation? -3x + 4y = -12

hartnn (hartnn):

To get x-intercept put y=0 in the equation -3x + 4y = -12 to get y-intercept put x=0 in the equation -3x + 4y = -12

OpenStudy (anonymous):

now im confused .-.

hartnn (hartnn):

put y=0 in -3x + 4y = -12 what you get ?

OpenStudy (anonymous):

-3x = -12 ?

hartnn (hartnn):

yes, so x=... ?

OpenStudy (anonymous):

4

hartnn (hartnn):

yes, so thats your x-intercept! (4,0) wasn't it easy ? :)

hartnn (hartnn):

now find y-intercpet by putting x=0

hartnn (hartnn):

4y=-12 y=-3 y-intercept = (0,-3) got it ?

OpenStudy (anonymous):

yesh. thank youu